Find a and b such that -3ai-(-1 - i) b =3a - 2bi
1
Expert's answer
2021-07-20T17:04:20-0400

Let us find "a" and "b" such that "-3ai-(-1 - i) b =3a - 2bi." It follows that "-3ai+b+bi =3a - 2bi," and hence "b+(b-3a)i =3a - 2bi". Taking into account that two complex numbers are equal if and only if their real parts are equal and their imaginary parts are equal, we conclude that  "b=3a, b-3a =- 2b." Then "b=3a" and "3b=3a." We conclude that "b=3b," which is equivalent to "2b=0," and hence "b=0". It follows that "a=b=0."


Answer"a=0,b=0."